OpenStudy (anonymous):

The one-to-one functions G and H are defined as follows. g={(-9,8),(-3,-5),(6,0),(8,1)} h(x)=3x+2 find the following: g^-1(8)= h^-1(x)= (hoh^1)(5)=

OpenStudy (zzr0ck3r):

g^-1 = {(8,-9),(-5,-3),(0,6),(1,8)} so g^-1(8) = -9 h^-1(x) = (x-2)/3 (y=3x+2 implies (y-2)/3 = x, then switch the x and y) (h dot h^-1)(x) = x so (h dot h^-1)(5) = 5

OpenStudy (anonymous):

thank you:)